APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Saturday October 31, 2026 (SJD 61345) ---- Observing Team ---- Kaike Pan ---- Observing Plan ---- Sat 31 Oct 2026 61345 0.57 01:07 01:22 12:12 21:00 07:51 --- dark until 05:34 ( 4.2 h), then bright ( 6.9 h, w/ 0.2 h for 12 deg design) ---- Observing Summary ---- Unfortunately, the entire night was lost to bad weather: high humidity, rain, fog. ---- Observing Log ---- Afternoon: ========== No adjustment made on BOSS collimation. 23:30:24 Hartmann collimate output sp1: offset: r= -659 (In Focus); b= -2.2 (In Focus) sp1: pred. move: spAverageMove= 24 sp1: pred. resid: r= -683, b= -2.1, txt= OK, spTemp = 7.2 BOSS cal, APOGEE evening cal, fvc dark, GFA bias, dark are taken, checked. The telescope, other systems and actors are checked, and confirmed work well. Robots are exercised, work as expected. No evening ventilation due to high humidity. Night: ====== Delayed open due to bad weather: rain, high humidity. Never open due to bad weather. ---- Mirror Numbers ---- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A: -5.4290 -5.4260 Axial B: -5.0390 -5.0390 Axial C: -5.2800 -5.2780 Trans D: -0.8760 -0.8710 Lateral E: 0.3240 0.3245 Lateral F: 0.5335 0.5335 GALILS Commanded: 11450 -1050 5900 26550 -14400 -13850 Actual: 11450 -1050 5900 26550 -14400 -13850 SETMIR VALUES PrimDesOrient: 0.0 -43.05 -8.77 -751.82 -556.2 PrimOrient: -1.06 -43.05 -8.77 -751.82 -556.2 SECONDARY: ---------- Focus: 0.0000 Air temp: 0.00 Altitude: 30.0 MIGS TONIGHT NOMINAL Axial A: -7.9560 -7.9570 Axial B: -6.1110 -6.1140 Axial C: -6.0750 -6.0790 Trans D: 0.2000 0.2050 GALILS Commanded: 363450 385700 345450 -1500 1850 Actual: 322072 350065 299539 -1500 1850 SETMIR VALUES SecDesOrient: 257.0 14.2 5.9 -72.0 -12.76 SecOrient: 257.04 14.24 5.85 -72.38 -13.22 ---- Telescope Status ---- N/A Started ventilation N/A encl25m : open N/A encl25m : closed No interlocks bypassed Status at: 12:00 Telescope stowed at: 121.0, 30.0, 0.0 mount Instrument mounted: FPS Counterweights at: LN2 autofill systems: Connected and turned on 180L LN2 dewar scales: SP1 192 lbs, 23.6 psi ---- Software Used ---- STUI : 1.7.21 alerts : 2.3.0 boss : 6.0.5a1 -boss.specMec : 3.0.4 apo : 4.1.0 fliswarm : 0.7.0 jaeger : 1.11.2a1 kaiju : 1.4.1 coordio : 1.18.1 fps_calibrati : 2025.12.04+3e7e2d8 cherno : 1.2.2 astrometrynet : 0.97 hal : 1.4.1 hartmann : 2.0.2 hub : 5.0.1a0 mcp : 2.3.1 -mcp.plc : 35 -mcp.azFiduci : v1_154 -mcp.altFiduc : v1_154 -mcp.rotFiduc : v1_154 tcc : 1.4.0 apogee : not availble apogeecal : not availble apogeeql : 2.0.0a0 STUIAdditions : FAILED Boson : 0.7.0 sdss-obstools : 3.9.3 TPM : 1.2-31-gd7eb MCP : 2.3.1 PLC : V35 Fiducials : v1_154 idlspec2d : 6.3.0dev Kronos : 2.1.0a roboscheduler : 2.1.0 ---- Time Tracking ---- *=*=*= Observatory: APO *=*=*= MJD: 61345 Fill night time in Hours below *=*=*= Total bright time : 6.9 *=*=*= Bright time, queued observations : 0 *=*=*= Bright time, commissioning/engineering: 0 *=*=*= Bright time lost to weather : 6.9 *=*=*= Bright time lost to technical issues : 0 *=*=*= Total dark time : 4.2 *=*=*= Dark time, queued observations : 0 *=*=*= Dark time, commissioning/engineering : 0 *=*=*= Dark time lost to weather : 4.2 *=*=*= Dark time lost to technical issues : 0